# Controlling Counterfactual Harm in Decision Support Systems Based on Prediction Sets

### v1

@@ 7 Discussion and limitations
Assumptions. We have assumed that the data samples and the expert predictions are drawn i.i.d. from a fixed distribution and the calibration set contains samples with noiseless ground truth labels and expert label predictions. It would be very interesting to extend our framework to allow for distribution shifts and label noise. Moreover, under the interventional monotonicity assumption, we have empirically observed that the gap between our lower- and upper-bounds is often large. Therefore, it would be useful to identify other natural assumptions under which this gap is smaller.
@@ 8 Conclusions
Appendix A Proofs
A.1 Proof of Proposition 1
@@
$=\mathbbm{1}\{\hat{y}=y\wedge y\notin\mathcal{C}_{\lambda}(x)\}.$
A.2 Proof of Proposition 2
@@
$\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[h_{\lambda}(x,y,\hat{Y}^{\prime})\right]$ $\leq\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{Y}^{\prime},X=x,Y=y}}\left[\mathbbm{1}\{\hat{Y}\neq y\}\right]\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}\right]$
$+\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}^{\prime}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}\right]$
$=\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{Y}^{\prime},X=x,Y=y}}\left[\mathbbm{1}\{\hat{Y}\neq y\}\right]\right]\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}$
@@
Using the above result, Eq. A.2 becomes
@@
$\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[h_{\lambda}(X,Y,\hat{Y})\right]\leq\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\}\cdot\mathbbm{1}\{Y\in\mathcal{C}_{\lambda}(X)\}+\mathbbm{1}\{\hat{Y}=Y\}\cdot\mathbbm{1}\{Y\notin\mathcal{C}_{\lambda}(X)\}\right],$
@@
$\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[h_{\lambda}(X,Y,\hat{Y})\right]\leq\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\wedge Y\in\mathcal{C}_{\lambda}(X)\}+\mathbbm{1}\{\hat{Y}=Y\wedge Y\notin\mathcal{C}_{\lambda}(X)\}\right].$
@@
$h_{\lambda}(x,y,\hat{y})$ $=\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{y},X=x,Y=y}}[\mathbbm{1}\{\hat{Y}\neq y\}]\cdot\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}+\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}$
@@ A.5 Additional Lemmas
$h_{\lambda}(x,y,\hat{y})=\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{y},X=x,Y=y}}[\mathbbm{1}\{\hat{Y}\neq y\}]\cdot\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}+\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}.$ (19)
@@
$\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,X=x,Y=y,\hat{Y}=\hat{Y}^{\prime}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\}\right]\right]=\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)}}[\mathbbm{1}\{\hat{Y}\neq Y\}\,|\,X=x,Y=y].$ (20)
@@
Proof. Let $\lambda,\lambda^{\prime}\in\mathcal{L}$ such that $\lambda\leq\lambda^{\prime}$. Then it will always hold that $\mathcal{C}_{\lambda}(x)\subseteq\mathcal{C}_{\lambda^{\prime}}(x)$. We distinguish the following cases for $\mathcal{C}_{\lambda}(x),\mathcal{C}_{\lambda^{\prime}}(x)$:
@@
Appendix B Additional experimental details
@@
Mixture of multinomial logit models. Multinomial logit models (MNLs) [49] are among the most popular discrete choice models used to predict the probability that a human chooses an alternative, given a set of alternatives within a particular context [50]. Straitouri et al. [18] used MNLs to predict the probability that a human expert predicts a label (e.g.the ground truth label), given a prediction set, while considering as context the ground truth label and the level of difficulty of the image. To this end, for images with the same level of difficulty, they estimate the confusion matrix of the expert predictions on their own and use it to parameterize an MNL model. Therefore, their mixture of MNLs comprises of a different MNL model for each level of difficulty. To distinguish images based on their level of difficulty, they consider different quantiles of the experts’ average accuracy values of all images. In our experiments, we consider two levels of difficulty, where in the first level we consider images for which the average accuracy of experts is smaller than the $0.5$ quantile, and consider the rest of the images in the second level.
@@
Appendix C Additional experimental results on counterfactual monotonicity
In this section, we show that the average accuracy $A(\lambda)$ achieved by a human participant using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ for $\alpha\in\{0.01,0.05\}$ on the strata of images with $\omega=80$ in Figure 4 and with $\omega=95$ in Figure 5. The meaning and coloring of each point are the same as in Figure 2. The results on both strata of images are consistent with the results on the strata of images with $\omega=110$, thus, supporting similar findings—the decision support systems $\mathcal{C}_{\lambda}$ always cause some amount of counterfactual harm, our framework successfully identifies harm-controlling sets and there is a trade-off between accuracy and counterfactual harm. However, for the strata with $\omega\in\{80,95\}$, where the classification task has lower difficulty compared to the strata with $\omega=110$, we observe that the trade-off between accuracy and counterfactual is minimal.
@@ Appendix C Additional experimental results on counterfactual monotonicity
Figure 4: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=80$. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.891$ (VGG19), $0.892$ (DenseNet161), $0.802$ (GoogleNet), $0.804$ (ResNet152), and $0.784$ (AlexNet). The average accuracy of human experts on their own is $0.9$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
@@
Figure 5: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=95$. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.88$ (VGG19), $0.868$ (DenseNet161), $0.775$ (GoogleNet), $0.773$ (ResNet152), and $0.745$ (AlexNet). The average accuracy of human experts on their own is $0.86$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
Appendix D Average accuracy vs. prediction set size

### v2

@@ 7 Discussion and limitations
Assumptions. We have assumed that the data samples and the expert predictions are drawn i.i.d. from a fixed distribution and the calibration set contains samples with noiseless ground truth labels and expert label predictions. It would be very interesting to extend our framework to allow for distribution shifts and label noise. Moreover, we have considered prediction sets constructed with a fixed user-specified threshold value $\lambda$. In light of recent work by Gibbs et al. [59], it would be interesting to extend our framework to allow for threshold values that depend on the data samples. Furthermore, in our definition of counterfactual harm we have treated all inaccurate predictions as equally harmful. Expanding our definition of harm to weigh different inaccurate predictions according to their consequences would be a very interesting avenue for future work. In addition, under the interventional monotonicity assumption, we have empirically observed that the gap between our lower- and upper-bounds is often large. Therefore, it would be useful to identify other natural assumptions under which this gap is smaller.
@@ 8 Conclusions
Appendix A Comparison to Pearl’s notation
Using Pearl’s notation [50], our definition of counterfactual harm reads as follows:
$h_{\lambda}(x,y,\hat{y})=\mathbb{E}[\max\{0,\mathbbm{1}\{\hat{y}=y\}-\mathbbm{1}\{\hat{Y}_{\lambda}=y\}\}|X=x,Y=y,\hat{Y}_{1}=\hat{y}],$
where the subindices in $\hat{Y}_{\lambda}$ and $\hat{Y}_{1}$ denote (hard) interventions on the exogenous variable $\Lambda$, the random variables $x,y,\hat{y}\sim P(X,Y,\hat{Y}_{1})$, and the expectation is over the prediction $\hat{Y}_{\lambda}$ made by the expert using the system $\mathcal{C}_{\lambda}$. Using the above notation, the definition does not explicitly specify the (counterfactual) distribution used in the expectation, in contrast to the definition using the notation by Peters et al. [51], which explicitly specifies the counterfactual distribution used in the expectation, as restated below for clarity.
$h_{\lambda}(x,y,\hat{y})=\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{y},X=x,Y=y}(\hat{Y})}[\max\{0,\mathbbm{1}\{\hat{y}=y\}-\mathbbm{1}\{\hat{Y}=y\}\}].$
Appendix B Proofs
B.1 Proof of Proposition 1
@@
$+\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{y},X=x,Y=y}}[\max\{0,1-0\}]\cdot\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}~{}$ (12)
@@
$=\mathbbm{1}\{\hat{y}=y\wedge y\notin\mathcal{C}_{\lambda}(x)\},$
where, in Eq. 12, we used the counterfactual monotonicity property by setting $\mathbbm{1}\{\hat{Y}=y\}=1$ if $y\in\mathcal{C}_{\lambda}(x)$.
B.2 Proof of Proposition 2
@@
$\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[h_{\lambda}(x,y,\hat{Y}^{\prime})\right]$ $\leq\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{Y}^{\prime},X=x,Y=y}}\left[\mathbbm{1}\{\hat{Y}\neq y\}\right]\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}\right]$ (13)
$+\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}^{\prime}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}\right]$ (14)
$=\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{Y}^{\prime},X=x,Y=y}}\left[\mathbbm{1}\{\hat{Y}\neq y\}\right]\right]\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}$ (15)
@@
Using the above result, Eq. B.2 becomes
@@
$\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[h_{\lambda}(X,Y,\hat{Y})\right]\\ \leq\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\}\cdot\mathbbm{1}\{Y\in\mathcal{C}_{\lambda}(X)\}+\mathbbm{1}\{\hat{Y}=Y\}\cdot\mathbbm{1}\{Y\notin\mathcal{C}_{\lambda}(X)\}\right],$
@@
$\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[h_{\lambda}(X,Y,\hat{Y})\right]\\ \leq\mathbb{E}_{X,Y,\hat{Y}\sim P^{\mathcal{M}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\wedge Y\in\mathcal{C}_{\lambda}(X)\}+\mathbbm{1}\{\hat{Y}=Y\wedge Y\notin\mathcal{C}_{\lambda}(X)\}\right].$
@@ B.5 Additional Lemmas
$h_{\lambda}(x,y,\hat{y})=\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,\hat{Y}=\hat{y},X=x,Y=y}}[\mathbbm{1}\{\hat{Y}\neq y\}]\cdot\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\in\mathcal{C}_{\lambda}(x)\}\\ +\mathbbm{1}\{\hat{y}=y\}\cdot\mathbbm{1}\{y\notin\mathcal{C}_{\lambda}(x)\}.$ (23)
@@
$\mathbb{E}_{\hat{Y}^{\prime}\sim P^{\mathcal{M}}}\left[\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)\,|\,X=x,Y=y,\hat{Y}=\hat{Y}^{\prime}}}\left[\mathbbm{1}\{\hat{Y}\neq Y\}\right]\right]\\ =\mathbb{E}_{\hat{Y}\sim P^{\mathcal{M};\text{do}(\Lambda=\lambda)}}[\mathbbm{1}\{\hat{Y}\neq Y\}\,|\,X=x,Y=y].$ (24)
@@
Proof.
Let $\lambda,\lambda^{\prime}\in\mathcal{L}$ such that $\lambda\leq\lambda^{\prime}$. Then it will always hold that $\mathcal{C}_{\lambda}(x)\subseteq\mathcal{C}_{\lambda^{\prime}}(x)$. We distinguish the following cases for $\mathcal{C}_{\lambda}(x),\mathcal{C}_{\lambda^{\prime}}(x)$:
@@
Appendix C Additional experimental details
@@
Mixture of multinomial logit models. Multinomial logit models (MNLs) [60] are among the most popular discrete choice models used to predict the probability that a human chooses an alternative, given a set of alternatives within a particular context [61]. Straitouri et al. [17] used MNLs to predict the probability that a human expert predicts a label (e.g., the ground truth label), given a prediction set, while considering as context the ground truth label and the level of difficulty of the image. To this end, for images with the same level of difficulty, they estimate the confusion matrix of the expert predictions on their own and use it to parameterize an MNL model. Therefore, their mixture of MNLs comprises of a different MNL model for each level of difficulty. To distinguish images based on their level of difficulty, they consider different quantiles of the experts’ average accuracy values of all images. In our experiments, we consider two levels of difficulty, where in the first level we consider images for which the average accuracy of experts is smaller than the $0.5$ quantile, and consider the rest of the images in the second level.
@@
Appendix D Additional experimental results on counterfactual monotonicity
In this section, we evaluate our methodology on additional strata of images and on larger calibration sets, and study the relationship between average counterfactual harm, average prediction set size and empirical coverage.
Other strata of images. Figures 4 and 5 show the average accuracy $A(\lambda)$ achieved by a human expert using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ for $\alpha\in\{0.01,0.05\}$ on the strata of images with $\omega=80$ and $\omega=95$, respectively. The meaning and coloring of each point are the same as in Figure 2. The results on both strata of images are consistent with the results on the stratum of images with $\omega=110$—the decision support systems $\mathcal{C}_{\lambda}$ always cause some amount of counterfactual harm, our framework successfully identifies harm-controlling sets and there is a trade-off between accuracy and counterfactual harm. However, for the strata with $\omega\in\{80,95\}$, where the classification task has lower difficulty compared to the stratum with $\omega=110$, we observe that the trade-off between accuracy and counterfactual is minimal.
@@ Appendix D Additional experimental results on counterfactual monotonicity
Figure 4: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=80$. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.891$ (VGG19), $0.892$ (DenseNet161), $0.802$ (GoogleNet), $0.804$ (ResNet152), and $0.784$ (AlexNet). The average accuracy achieved by human experts on their own is $0.9$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
@@
Figure 5: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=95$. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.88$ (VGG19), $0.868$ (DenseNet161), $0.775$ (GoogleNet), $0.773$ (ResNet152), and $0.745$ (AlexNet). The average accuracy achieved by human experts on their own is $0.86$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
Larger calibration sets. Figure 6 shows the average accuracy $A(\lambda)$ achieved by a human expert using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ for $\alpha=0.01$ on the stratum of images with $\omega=110$. The results show that the larger the calibration set, the more often $\lambda$ values with average counterfactual harm $H(\lambda)$ close to $\alpha$ [^12] Here, note that for $\alpha=0.01$, these $\lambda$ values happen to achieve the highest average accuracy $A(\lambda)$. are included in the harm-controlling set $\Lambda(\alpha)$. This is because, the larger the number of samples in the calibration set ($n$), the lower the estimation error in the empirical estimate of the average counterfactual harm $\hat{H}_{n}$. We find qualitatively similar results for the other strata of images.
(a) Calibration samples $n=360$ $(30\%)$
(b) Calibration samples $n=600$ $(50\%)$
Figure 6: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=110$ and different sizes of the calibration set. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.846$ (VGG19), $0.830$ (DenseNet), $0.722$ (GoogleNet), $0.727$ (ResNet152), and $0.691$ (AlexNet). The average accuracy achieved by the simulated human experts on their own is $0.771$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals are represented using shaded areas and always have width below $0.02$.
Average prediction set-size and empirical coverage. Figure 7(a) shows the average prediction set-size of each decision support system $\mathcal{C}_{\lambda}$ against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ on the stratum of images with $\omega=110$. Figure 7(b) shows the empirical coverage, i.e., the fraction of samples $x$ in the test set, for which ground truth label $y\in\mathcal{C}_{\lambda}(x)$, for each decision support system $\mathcal{C}_{\lambda}$ against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ on the same stratum of images. The results show that systems with higher coverage construct larger prediction sets in average and cause less counterfactual harm. We find similar results for the other strata of images.
(a) Average Set-Size
(b) Empirical Coverage
Figure 7: Average set size and empirical coverage against the average counterfactual harm for images with $\omega=110$. Each point corresponds to a $\lambda$ value from 00 to $1$ with step $0.001$. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies as in Figure 6 above. The average accuracy achieved by the simulated human experts on their own is $0.771$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals are represented using shaded areas and always have width below $0.025$.
Appendix E Experimental results on counterfactual monotonicity with a more complex set-valued predictor
In this section, we experiment with a system that constructs the prediction set $\mathcal{C}(x)$ using a more complex set-valued predictor [49]. This set-valued predictor first computes the following nonconformity score for each label value $y\in\mathcal{Y}$ using the softmax output of a pre-trained classifier $m_{y}(x)\in[0,1]$:
$s(y,x,u;w)=\left\{\begin{array}[]{cc}u\cdot m_{y_{(1)}}(x),&\text{if }y=y_{(1)}\\ m_{y_{(1)}}+(i-2+u)\cdot w,&\text{if }y=y_{(i)}\end{array}\right.,$ (28)
where $y_{(i)}$ denotes the label value with the $i$-th largest classifier output, $u\sim U(0,1)$ is a uniform random variable and $w$ is a hyperparameter representing the weight of the ranking of the label value. Then, given a user-specified threshold $\lambda\in\mathbb{R}$, it constructs the prediction sets $\mathcal{C}(x)=\mathcal{C}_{\lambda}(x)$ as follows:
$\mathcal{C}_{\lambda}(x)=\{y_{(L-i)}\}_{i=0}^{k-1},\text{ with }k=1+\sum_{j=1}^{L-1}\mathbbm{1}\{s(y_{(j)},x,u;w)\leq\lambda\},$ (29)
where, with a slight abuse of notation, $y_{(i)}$ is the label value with the $i$-th largest $s(y,x,u;w)$ value. Here, we should note that may exist data samples with $s(y,x,u;w)>1$. Therefore, considering $\lambda\in[0,1]$ may not include every set-valued predictor given by Eq. 29. For this reason, in what follows, we consider $\lambda\in[0,\max s(y,x,u,w)]$, where the maximum is essentially over the classifier output $m_{y}(x)$, the uniform random variable $u$ and the hyperparameter $w$.
Experimental setup. We use the ImageNet16H dataset and follow the same setup as described in Section 6, with only difference that we split the images into a calibration set (10%), which we use to find $\Lambda(\alpha)$, a validation set $(10\%)$, which we use to select the value of the hyperparameter $w$ following a similar procedure as in Huang et al. [49], and a test set (80%). In each iteration of the experiment, given a $\lambda$ value, we select the value of $w\in\{0.02,0.05,0.1,0.15,0.2,0.25,0.3,0.35\}$ for which the set-valued predictor achieves the smallest average prediction set-size over the data samples in the validation set.
Results. The meaning and coloring of each point are the same as in Figure 2. The results on all strata of images are consistent with the results using the set-valued predictor given by Eq. 1—the decision support systems $\mathcal{C}_{\lambda}$ always cause some amount of counterfactual harm, our framework successfully identifies harm-controlling sets and there is a trade-off between accuracy and counterfactual harm.
Figure 11(a) shows the average prediction set-size of each decision support system $\mathcal{C}_{\lambda}$ against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ on the stratum of images with $\omega=110$. Figure 11(b) shows the empirical coverage for each decision support system $\mathcal{C}_{\lambda}$ against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ on the stratum of images with $\omega=110$. Our results are consistent with the results for the set-valued predictor given by Eq. 1 in Figure 7—systems $\mathcal{C}_{\lambda}$ that achieve higher coverage, construct larger prediction sets on average and cause less harm.
(a) $\alpha=0.01$
(b) $\alpha=0.05$
Figure 8: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=80$. Each point corresponds to a $\lambda$ value from 00 to $6.25$ with step $0.00625$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.891$ (VGG19), $0.892$ (DenseNet161), $0.802$ (GoogleNet), $0.804$ (ResNet152), and $0.784$ (AlexNet). The average accuracy of human experts on their own is $0.9$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
(a) $\alpha=0.01$
(b) $\alpha=0.05$
Figure 9: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=95$. Each point corresponds to a $\lambda$ value from 00 to $6.25$ with step $0.00625$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.88$ (VGG19), $0.868$ (DenseNet161), $0.775$ (GoogleNet), $0.773$ (ResNet152), and $0.745$ (AlexNet). The average accuracy of human experts on their own is $0.86$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals have width always below $0.02$ and are represented using shaded areas.
(a) $\alpha=0.01$
(b) $\alpha=0.05$
Figure 10: Average accuracy estimated by the mixture of MNLs against the average counterfactual harm for images with $\omega=110$. Each point corresponds to a $\lambda$ value from 00 to $6.25$ with step $0.00625$ and the coloring indicates the relative frequency with which each $\lambda$ value is in $\Lambda(\alpha)$ across random samplings of the calibration set. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies $0.846$ (VGG19), $0.830$ (DenseNet), $0.722$ (GoogleNet), $0.727$ (ResNet152), and $0.691$ (AlexNet). The average accuracy achieved by the simulated human experts on their own is $0.771$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals are represented using shaded areas and always have width below $0.02$.
(a) Average Set-Size
(b) Empirical Coverage
Figure 11: Average set size and empirical coverage against the average counterfactual harm for images with $\omega=110$. Each point corresponds to a $\lambda$ value from 00 to $6.25$ with step $0.00625$. Each row corresponds to decision support systems $\mathcal{C}_{\lambda}$ with a different pre-trained classifier with average accuracies as in Figure 10 above. The average accuracy achieved by the simulated human experts on their own is $0.771$. The results are averaged across $50$ random samplings of the test and calibration set. In both panels, $95\%$ confidence intervals are represented using shaded areas and always have width below $0.025$.
Appendix F Average accuracy vs. prediction set size